A source emits electromagnetic waves of wavelength 3m. One beam reaches the observer
directly and other after reflection from a water surface, travelling 1.5m extra distance
and with intensity reduced to 1/4 as compared to intensity due to the direct beam alone.
The resultant intensity will be
1) (1/4) fold 2) (3/4) fold 3) (5/4) fold 4) (9/4) fold
1
Expert's answer
2019-05-03T09:00:02-0400

The resultant amplitude is equal to


"\sqrt{I}+\sqrt{\frac {1} {4}I}=\frac {3} {2} \sqrt{I}"

The resultant intensity is equal to


"({\frac {3} {2} \sqrt{I}})^2=\frac {9} {4} \sqrt{I}"

Answer:

4) (9/4) fold